OpenStudy (anonymous):

IM SOO CONFUSED!!! 4x-3y+5z=43 2x+y=9 3y-22=-9 please show me how to do these

OpenStudy (saifoo.khan):

Confusions are good.

OpenStudy (anonymous):

really?

OpenStudy (saifoo.khan):

they really makes a person cinfuse.

OpenStudy (saifoo.khan):

confuse*

OpenStudy (anonymous):

hmm.. i will have to thnk bout that

OpenStudy (anonymous):

comon seafood i need help!!

OpenStudy (anonymous):

HELLP!!! (and no im not gonna call 911 seafood)

OpenStudy (anonymous):

hellp!!!!!

OpenStudy (anonymous):

i jus need sumbody to explain it for me!!!

OpenStudy (anonymous):

hi ayesha

OpenStudy (whynot):

hi GoodMan

OpenStudy (whynot):

i think you can find the value of y in the last eqation and plug that in second to get x and plu both in the first to get z

OpenStudy (anonymous):

so i gotta find the value of y first... AND NO ITS NOT INCORRECT SEAFOOD!!

OpenStudy (whynot):

yes find y . Plug that in the second to find x and finally plug both in first

OpenStudy (anonymous):

WAIT...the last equation was 3y-2z=-9

OpenStudy (saifoo.khan):

So? it's right.

OpenStudy (anonymous):

wadda ya mean its right

OpenStudy (saifoo.khan):

What the heck are your teachers doing back @ skol?

OpenStudy (anonymous):

seriously i dunt know wat ur taking bout seafood

OpenStudy (saifoo.khan):

why dont u learn that at school? stop sleeping in the class.

OpenStudy (anonymous):

I DUNT SLEEP IN CLASS!!!!

OpenStudy (saifoo.khan):

THen? stop dreaming in the class.

OpenStudy (saifoo.khan):

whynot, explain her.

OpenStudy (anonymous):

OMG....can u plz jus answer the question step by step...im kinda stuck

OpenStudy (whynot):

find z from last one. Find x from secong plug both in the first and you will get y that way

OpenStudy (anonymous):

hmm..lemme try

OpenStudy (whynot):

ok

OpenStudy (anonymous):

wait..so find y first by substituion?

OpenStudy (saifoo.khan):

let me explain this simple crap.

OpenStudy (anonymous):

ahh...jus be quiet seafood!! u need help too time to time! NO BODY IS PERFECT!!! and plz let tat stick into ur mind

OpenStudy (anonymous):

THANK YOU AYESHA!!!

OpenStudy (saifoo.khan):

LOL, i read it... Anyways, i was just trying to kid around with someoneelseman

OpenStudy (whynot):

Hey Saifoo, plz dont be rude.We are all learning:)

OpenStudy (anonymous):

...now here comes another insult from seafood....

OpenStudy (saifoo.khan):

4x-3y+5z=43 2x+y=9 3y-22=-9 Let's take equation 2 first. 2x + y = 9 Solve this for any variable you like, let;s do it for y. y = 9 - 2x. Now you have the value y. Insert that y into equation 3, instead of y in it. 3y-22=-9 3(9-2x) - 22 = -9 Solve it to get the value for x.

OpenStudy (anonymous):

nm

OpenStudy (saifoo.khan):

Haha, No.. here comes another "HELP" from seefood.

OpenStudy (anonymous):

*seafood

OpenStudy (whynot):

lol

OpenStudy (saifoo.khan):

lol, that one will work as well.

OpenStudy (saifoo.khan):

Stop chating around and try to learn that question!

OpenStudy (anonymous):

thank you seafood..(u started it)

OpenStudy (saifoo.khan):

No, you did.. Always girls start bad things! :D

OpenStudy (whynot):

Not True!!

OpenStudy (saifoo.khan):

It is true.

OpenStudy (anonymous):

gurls are considered smarter than guys studies show...so ya noe wat? u started it

OpenStudy (saifoo.khan):

girls*

OpenStudy (saifoo.khan):

know*

OpenStudy (anonymous):

that was no typo seafood

OpenStudy (saifoo.khan):

And your second last comment was a joke, i know that.

OpenStudy (saifoo.khan):

whynot, how's dengue in Lahore? :O

OpenStudy (anonymous):

wat comment?

OpenStudy (saifoo.khan):

"gurls are considered smarter than guys studies show...so ya noe wat? u started it "

OpenStudy (anonymous):

its is sooo true...look it up!

OpenStudy (saifoo.khan):

where is it? i can't see it.

OpenStudy (whynot):

It is true!!1 and my lovely Lahore is doing very good Alhamdulillah

OpenStudy (saifoo.khan):

i know it's true that girls always start wrong things. :P very good with dengue? :O

OpenStudy (anonymous):

2 against 1! u lose seafood..GURLS ARE SMARTER THAN GUYS!!!

OpenStudy (whynot):

No ! it is getting better now

OpenStudy (whynot):

yup Girls are and we dont need to prove that:)

OpenStudy (anonymous):

tat also proves seafood as a LOSER!!

OpenStudy (saifoo.khan):

lol^

OpenStudy (whynot):

It is a universal truth

OpenStudy (anonymous):

God sent us girls to expand the future...and we are doing a perty good job on that!

OpenStudy (saifoo.khan):

Goodman is really a badman, not intrested in learning his problem instead of it.. Joking around like a joker. :D

OpenStudy (saifoo.khan):

Haha, Another joke by badman.

OpenStudy (anonymous):

was that a poem?

OpenStudy (saifoo.khan):

I never said it is a poen

OpenStudy (saifoo.khan):

poem*

OpenStudy (whynot):

well, anyway we are all good in our own way.

OpenStudy (anonymous):

yep...well tanx for the help ayesha!!!

OpenStudy (whynot):

with just an exception that girls are a little bit more smart

OpenStudy (whynot):

No problem:)

OpenStudy (saifoo.khan):

Lame Jokes. :O

OpenStudy (anonymous):

believe wat u believe..but deep down the truth is tare and u jus dunt wanna accept it !

OpenStudy (saifoo.khan):

again Joke.

OpenStudy (whynot):

lol

OpenStudy (anonymous):

its not a joke wen ayesha says it too!

OpenStudy (saifoo.khan):

lol, ayesha is laughing -that means it is a JOKE. hahaha

OpenStudy (anonymous):

well ayesha wat do u thnk?

OpenStudy (whynot):

i am laughing at your comment saifoo u seem to make everything into a joke

OpenStudy (saifoo.khan):

LOL, it is ajoke.

OpenStudy (anonymous):

HAHAHA!! LOSER!!! U GOT SERVED!!!

OpenStudy (whynot):

hahaha! hay goodman lets saifoo loose .It is ok if he wants to

OpenStudy (saifoo.khan):

LOL, who said i lost. I'm not a loser. Infact you are a Joker . :P

OpenStudy (whynot):

it is not your fault.You are told that .It is ok.No one is going to hurt u

OpenStudy (anonymous):

u sooo lost seafood!!!

OpenStudy (saifoo.khan):

No one can even hurt me! :D

OpenStudy (whynot):

yeah.Even an ant can bite u!!!

OpenStudy (saifoo.khan):

lol, you always joke around, that;s why u dont get anything wht your teacher teaches u.

OpenStudy (anonymous):

LOL

OpenStudy (saifoo.khan):

lol, No it dosn't. we dont have many in Khi

OpenStudy (whynot):

hy! dont be mean now.You know we are joking

OpenStudy (whynot):

We appreciate that you help students around here:)

OpenStudy (saifoo.khan):

hy! dont be mean now.You know I'm joking as well

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

thanks for the appreciation.

OpenStudy (whynot):

:) ok gtg guys